OpenStudy (anonymous):

Let C denote the positively oriented boundary of the square whose sides lie along the lines x=(+/-)2 and y=(+/-)2. Evaluate the integral, over C of (e^(-z))/(z-((i*pi)/2))

OpenStudy (anonymous):

rewrite it again

OpenStudy (anonymous):

what are the lines

OpenStudy (anonymous):

ohh its +- 2 , man so stupid

OpenStudy (anonymous):

so the singularity is at \[\frac{i \pi}{2}\]

OpenStudy (anonymous):

and pi/2 is about 1.7 , so its inside the contour

OpenStudy (anonymous):

so all you have to do is sub \[\frac{i \pi }{2}\] into the function on top

OpenStudy (anonymous):

so you get -1

OpenStudy (anonymous):

and thats the answer

OpenStudy (anonymous):

complex analysis for the win

OpenStudy (anonymous):

I am told that the answer is pi?

OpenStudy (anonymous):

I mean 2pi

OpenStudy (anonymous):

ohh yeh , I forgot to multiply by the 2pi i fudge factor

OpenStudy (anonymous):

no its actually -2 pi i

OpenStudy (anonymous):

ur confusing me :|

OpenStudy (anonymous):

so I plug (i*Pi)/2 into e^(-z) and then multiply by 2*pi*i?

OpenStudy (anonymous):

yes, thats the cauchy integral formula isnt it

OpenStudy (anonymous):

yes, thank you!!!!!

OpenStudy (anonymous):

if you have \[\int\limits \frac{f(z)}{(z-a)^{n+1}} = 2 \pi i \frac{ f ^{(n)} (a) } {n!}\]

OpenStudy (anonymous):

so you differentite it zero times ( which does nothing ) , and divide by 0! (which is 1 )

OpenStudy (anonymous):

how do I mark this as a good answer?

OpenStudy (anonymous):

Thats the formula I was looking for. It really is that simple. However, that specific form only works for problems in the form. If you have multiple poles in the denominator it can make the problem different.

OpenStudy (anonymous):

I personally prefer the more general questions where you actually have to find the residues

OpenStudy (anonymous):

or the ones like (1/sinz) , where you come up with a general odd series and multiply both sides by sinz , compare coeffiecents etc

OpenStudy (anonymous):

but they are a bit longer.

OpenStudy (anonymous):

most singularities are simple poles though , can get the resides in like 30seconds by L hopitals rule

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so stupid lols

OpenStudy (anonymous):

well I have been on hollidays for like 3weeks

OpenStudy (anonymous):

\[e^{ \frac{-\pi i }{2}} = -i \]

OpenStudy (anonymous):

=\[2 \pi i \times -i = 2\pi \]

OpenStudy (anonymous):

so many mistakes, not really my day, lucky dont have a test today

OpenStudy (anonymous):

The Cauchy Integral Formula is \[f(s) = (i/(2 \pi i))\int\limits_{C}^{?}f(z) dz/(z - s).\]Here \[f(z) = e^{-z}, s = \pi i/2,\] \[f(s) = e^{-\pi i/2} = -i,\] so \[-i = \left( 1/(2 \pi i \right)\int\limits_{c}^{?}e^{-z}dz / (z - \pi i / 2),\] since pi * i/2 is in the interior of C. Hence,\[\int\limits_{C}^{?}e^{-z} dz/(z - \pi i/2) = 2 \pi.\]